# Potential and Limitation of High-Frequency Cores and Caches

### v1

@@ Abstract
This paper explores the potential of cryogenic computing and superconducting electronics as promising alternatives to traditional semiconductor devices. As semiconductor devices face challenges such as increased leakage currents and reduced performance at higher temperatures, these novel technologies offer high performance and low power computation. Cryogenic computing operates at ultra-low temperatures near 77 K, leading to lower leakage currents and improved electron mobility. On the other hand, superconducting electronics, operating near 0 K, allow electrons to flow without resistance, offering the potential for ultra-low-power, high-speed computation. This study presents a comprehensive performance modeling and analysis of these technologies and provides insights into their potential benefits and limitations. We implement models of in-order and out-of-order cores operating at high clock frequencies associated with superconducting electronics and cryogenic computing in gem5. We evaluate the performance of these components using workloads representative of real-world applications like NPB, SPEC CPU2006, and GAPBS. Our results show the potential speedups achievable by these components and the limitations posed by cache bandwidth. This work provides valuable insights into the performance implications and design trade-offs associated with cryogenic and superconducting technologies, laying the foundation for future research in this field using gem5. Our code and data are open-source and available on GitHub [^1] https://github.com/darchr/gem5-cryo-superconducting.
@@ I Introduction
Cryogenic computing, operating at ultra-low temperatures near 77 K, presents a promising avenue on high-speed and low-power computation [1]. Another technology, superconducting electronics, operating at near 0 K, offers the potential for ultra-low-power, high-speed computation, by allowing electrons to flow without resistance [3]. These technologies allow for individual components to operate at a higher clock frequency than typical CMOS devices.
In this paper, we study the full-system implications of running the processing core at ultra-high frequencies. Specifically, we introduce a modeling framework and perform and initial evaluation of out-of-order and in-order cores at high clock frequencies associated with superconducting electronics and cryogenic computing. We also wish to substantiate how these speedups will affect other components in the system, and how these other components should be designed to keep up with the faster components.
To evaluate the performance of cryogenic and superconducting computing technologies, we:
@@
    Use gem5 [4, 5] to implement models of in-order and out-of-order cores operating at high clock frequencies, typical of cryogenic and superconducting environments.
@@
    Choose workloads representative of real-world applications like NPB [6], SPEC CPU2006 [7], and GAPBS [8] to evaluate the performance of these components. These workloads provide a comprehensive understanding of how different types of applications respond to the ultra-fast processing capabilities offered by cryogenic and superconducting technologies.
@@
    Simulate these workloads in different combinations of the aforementioned components in a superconducting clock domain and in the cryogenic clock domain, while keeping the board and memory at their room temperature frequencies.
@@ II Background
In recent years, the pursuit of novel computing paradigms has led to a resurgence of interest in cryogenic components and superconducting circuits. These technologies promise unprecedented speedups by exploiting the unique properties of materials at extremely low temperatures.
While superconducting was first discovered in 1911 [9], the theoretical concepts behind it were not fully understood until the BCS theory of the 1950s [10]. Due to further work in concepts like the Josephson effect [11] and flux quantization [12], superconducting circuits have been understood to operate at high clock frequencies of up to 100 GHz [13]. Therefore, superconducting circuits enable ultra-fast signal processing with minimal energy dissipation [3].
Cryogenic components offer improved performance and energy efficiency by the reduced resistance at ultra-low temperatures. At cryogenic temperatures, semiconductor devices exhibit lower leakage currents and improved electron mobility, leading to improved performance. Therefore, cryogenic components also operate at a higher clock frequency, usually around 4 GHz [1].
Cryogenic computing has gained significant attention in recent years, with researchers exploring cryogenic components to enhance computing performance. Byun et al. introduced CryoCore, a cryogenic core based on the BOOM core, operating at 4.0 GHz with improved performance at cryogenic temperatures [1]. Min et al. proposed CryoCache, a cryogenic cache hierarchy leveraging these temperatures to enhance cache performance [14]. These components operate at ultra-low temperatures and offer substantial speedups compared to conventional semiconductor devices.
Our work represents a crucial step towards the practical implementation of superconducting and cryogenic components in computing systems. Notably, we:
@@
    Lay the foundation for gem5 as a modeling framework for this field of computing research, providing all the tools necessary for evaluating the performance of computing systems under cryogenic and superconducting conditions.
@@ III Methodology
For all the experiments, we used gem5 [5, 4] as the simulator. We leverage the easy modification of gem5 CPU and cache models to make our own custom components, which is essential due to the microarchitectural details of the cryogenic computing models available in the literature. We can also place the CPU, caches, and memory in different clock domains, which is essential for simulating cryogenic computing environments. The gem5 Resources artifacts [15] provides workloads that we can also use for our experiments. Hence, gem5 is the ideal choice for our experiments.
@@ III-A Core Microarchitecture
We modeled the cryogenic core and caches by building a model based on CryoCore by Byun et al. [1]. Their cryogenic core was based on BOOM [16], an out-of-order RISC-V core, and could run at 4.0 GHz. We used microarchitectural details provided in their paper to create a variant of the O3CPU in gem5. These microarchitectural details are enumerated in Table 3. For the pipeline values that were not specified, we went with the default values provided by gem5 for those variables. With regard to the latencies of the functional units and the branch predictor, we decided to go with the ones provided as part of the RISCVMatched prebuilt board [17] (based on the HiFive Unmatched and having a similar architecture to Rocket and BOOM [18]), since these values have been partially validated [19].
@@ III-D Experimental Setup
  • CryoCore and CryoCache (CryoAll) - Out-of-order CryoCore and CryoCache models in the cryogenic clock domain (4 GHz).
  • SuperCore and CryoCache (SuperCryo) - CryoCore model in the superconducting clock domain (100 GHz) and CryoCache model in the cryogenic clock domain (4 GHz).
  • SuperCore and SuperCache (SuperAll) - CryoCore and CryoCache models in the superconducting clock domain (100 GHz).
  • In-Order CryoCore and CryoCache (In-Order CryoAll) - In-order CryoCore model and CryoCache model in the cryogenic clock domain (4 GHz).
  • In-Order SuperCore and CryoCache (In-Order SuperCryo) - In-order CryoCore model in the superconducting clock domain (100 GHz) and CryoCache model in the cryogenic clock domain (4 GHz).
  • In-Order SuperCore and SuperCache (In-Order SuperAll) - In-order CryoCore and CryoCache models in the superconducting clock domain (100 GHz).
@@ IV Results
Research Questions 1. How does the performance of workloads vary with increasing clock frequency when subsystems are placed in a superconducting environment? 2. How is the performance of workloads affected by whether the core is out-of-order or in-order? 3. What new constraints must subsystems meet to leverage the improved performance resulting from increased clock frequency in a superconducting environment?
IV-A Workload Performance with Increasing Clock Frequency in Superconducting Environments
@@
Figure 2(a) and 2(b) show the speedup of the small workloads with respect to the out-of-order and in-order cores. Specifically, it compares the speedups achieved by placing the core and/or cache in a superconducting environment, for both out-of-order and in-order cores.
@@
At a high level, the takeaways from these results are that placing both core and cache in a superconducting environment significantly boosts performance, and that speedups of in-order superconducting components over their cryogenic counterparts are generally higher than the speedups of out-of-order superconducting components over their cryogenic counterparts. By characterizing the workloads that benefit the most from superconducting components, we also provide insights into the design trade-offs associated with these novel technologies.
@@
For out-of-order cores, Figure 2(a) shows the speedup of the small workloads in various out-of-order core configurations, with respect to the out-of-order CryoCore and CryoCache configuration. The bars represent different configurations for different small workloads. If only the core is placed in the superconducting environment, the performance improvement is not substantial, with the maximum speedup for small workloads being 2.7$\times$ for the cg workload over the baseline for an out-of-order core. Overall, the speedups for the small workloads range from 1.8$\times$ to 2.7$\times$ for the out-of-order SuperCore and CryoCache configuration, compared to the out-of-order CryoCore and CryoCache configuration. If both the core and the cache are placed in the superconducting environment, the performance improvement is more substantial for the out-of-order core. The maximum speedup for the small workloads for an out-of-order SuperCore and SuperCache configuration is 23.2$\times$ for the bt workload over the baseline out-of-order CryoCore and CryoCache configuration. Some workloads like lu and cg also receive speedups of more than 20$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration. However, some workloads like bfs and minisat do not receive substantial speedups, with the speedup being around 9.4$\times$ and 7.8$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration, respectively. Overall, the speedups for the small workloads range from 7.8$\times$ to 23.2$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration.
@@
Figure 2(c) shows the speedup of the large workloads in various out-of-order core configurations, with respect to the out-of-order CryoCore and CryoCache configuration. The bars represent different configurations for different SPEC2006 workloads. Similar to the small workloads, if only the core is placed in the superconducting environment, the performance improvement is not substantial, with the maximum speedup being 3.9$\times$ for the 400.perlbench workload over the baseline for an out-of-order core. Overall, the speedups for the large workloads range from 1.01$\times$ to 3.9$\times$, barring 456.hmmer, for the out-of-order SuperCore and CryoCache configuration, compared to the out-of-order CryoCore and CryoCache configuration. If both the core and the cache are placed in the superconducting environment, the performance improvement is more substantial for the out-of-order core. The maximum speedup for the large workloads for an out-of-order SuperCore and SuperCache configuration is 24.0$\times$ for the 456.hmmer workload over the baseline out-of-order CryoCore and CryoCache configuration. Some workloads like 435.gromacs and 401.bzip2 also receive speedups of more than 20$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration. However, most workloads do not receive substantial speedups, with the speedup being around 1.2$\times$ and 1.7$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration, respectively. Overall, the speedups for the large workloads range from 1.18$\times$ to 24.0$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration.
@@
For in-order cores, Figure 2(b) shows the speedup of the workloads in various in-order core configurations, with respect to the in-order CryoCore and CryoCache configuration. The bars represent different configurations for different small workloads. If only the core is placed in the superconducting environment, the performance improvement is not substantial, with the maximum speedup for the small workloads being 2.3$\times$ for the tc workload over the baseline for an in-order core. Overall, the speedups for the small workloads range from 1.4$\times$ to 2.3$\times$ for the in-order SuperCore and CryoCache configuration, compared to the in-order CryoCore and CryoCache configuration. If both the core and the cache are placed in the superconducting environment, the performance improvement is more substantial for both the in-order core. The maximum speedup for an in-order SuperCore and SuperCache configuration is 24.3$\times$ for the bt workload over the baseline in the in-order SuperCore and CryoCache configuration. Some workloads like lu and cg also receive speedups of more than 20$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration. However, some workloads like bfs and minisat do not receive substantial speedups, with the speedup being around 8.4$\times$ and 9.6$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration, respectively. Overall, the speedups for the small workloads range from 8.4$\times$ to 24.3$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration.
@@
Figure 2(d) shows the speedup of the large workloads in various in-order core configurations, with respect to the in-order CryoCore and CryoCache configuration. The bars represent different configurations for different SPEC2006 workloads. Similar to the small workloads, if only the core is placed in the superconducting environment, the performance improvement is not substantial, with the maximum speedup being 2.6$\times$ for the 433.milc workload over the baseline for an in-order core. Overall, the speedups for the large workloads range from 1.1$\times$ to 2.6$\times$ for the in-order SuperCore and CryoCache configuration, compared to the in-order CryoCore and CryoCache configuration. If both the core and the cache are placed in the superconducting environment, the performance improvement is more substantial for both the in-order core. The maximum speedup for the large workloads for an in-order SuperCore and SuperCache configuration is 24.5$\times$ for the 456.hmmer workload over the baseline in the in-order SuperCore and CryoCache configuration. Some workloads like 435.gromacs, 444.namd and 401.bzip2 also receive speedups of more than 20$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration. However, most workloads do not receive substantial speedups, with the speedup being around 1.4$\times$ and 2.6$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration, respectively. Overall, the speedups for the large workloads range from 1.4$\times$ to 24.5$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration.
@@
Therefore, the speedups vary significantly across different workloads, with similar workloads achieving similar speedups across different core architectures. While the speedups are consistent for smaller workloads, the variation in speedup is much wider when running real-world workloads. This variation suggests that even if we could build a conventional architecture in a superconducting environment, it is unlikely to benefit general-purpose workloads uniformly. Therefore, we need to either develop a new architecture specifically optimized for superconducting environments or apply superconducting technology to specialized accelerators where there is a small amount of memory traffic instead of general-purpose compute, which could leverage the low latency and high throughput benefits of superconductors more effectively.
@@
Takeaways The performance of workloads improved by placing components in the superconducting environment. The performance improvement was more substantial when both the core and the cache were in the superconducting environment. In-order cores were able to achieve a higher speedup than out-of-order cores when both the core and the cache were placed in the superconducting environment. The workloads that did not receive substantial speedups were bottlenecked by memory accesses. Therefore, memory systems would need to be redesigned to keep up with the faster core. The speedups vary significantly across different workloads, with similar workloads achieving similar speedups across different core architectures. While the speedups are consistent for smaller workloads, the variation in speedup is much wider when running real-world workloads. Therefore, we need to either develop a new architecture specifically optimized for superconducting environments or apply superconducting technology to specialized accelerators where there is a small amount of memory traffic instead of general-purpose compute.
@@
We compared the speedups of various configurations of out-of-order and in-order superconducting and cryogenic cores and caches, with respect to the out-of-order CryoCore and CryoCache configuration. At a high level, we observed that the speedups are more substantial for out-of-order configurations compared to in-order configurations.
@@
From these figures, we can take away that for small workloads, placing components in a superconducting environment improves performance for both out-of-order and in-order cores. The performance improvement is more substantial when both the core and cache are in the superconducting environment. Interestingly, in-order cores achieve higher speedups compared to out-of-order cores when both components are superconducting. This could be because in-order cores are simpler and have less instruction-level parallelism, which could benefit from the higher clock frequency.
The out-of-order SuperCore and CryoCache configuration performs better than or equal to both the in-order CryoCore and CryoCache configuration and the in-order SuperCore and SuperCache configuration, for both the small and large workloads. Out-of-order cores can schedule instructions dynamically, which could benefit more from the higher clock frequency compared to in-order cores, which have a fixed schedule of instructions. Thus, even when all components are not superconducting, the out-of-order cores can still achieve significant performance benefits compared to in-order cores.

### v2

@@ Abstract
This paper explores the potential of cryogenic semiconductor computing and superconductor electronics as promising alternatives to traditional semiconductor devices. As semiconductor devices face challenges such as increased leakage currents and reduced performance at higher temperatures, these novel technologies offer high performance and low power computation. Conventional semiconductor electronics operating at cryogenic temperatures (below -150° C or 123.15 K) can benefit from reduced leakage currents and improved electron mobility. On the other hand, superconductor electronics, operating below 10 K, allow electrons to flow without resistance, offering the potential for ultra-low-power, high-speed computation. This study presents a comprehensive performance modeling and analysis of these technologies and provides insights into their potential benefits and limitations. We implement models of in-order and out-of-order cores operating at high clock frequencies associated with superconductor electronics and cryogenic semiconductor computing in gem5. We evaluate the performance of these components using workloads representative of real-world applications like NPB, SPEC CPU2006, and GAPBS. Our results show the potential speedups achievable by these components and the limitations posed by cache bandwidth. This work provides valuable insights into the performance implications and design trade-offs associated with cryogenic and superconductor technologies, laying the foundation for future research in this field using gem5. Our code and data are open-source and available on GitHub [^1] https://github.com/darchr/gem5-cryo-superconducting.
@@ I Introduction
Conventional semiconductor electronics operating at cryogenic temperatures (below -150° C or 123.15 K), present a promising avenue on high-speed and low-power computation [1]. Another technology, superconductor electronics, operating below 10 K, offers the potential for ultra-low-power, high-speed computation, by allowing electrons to flow without resistance [3]. These technologies allow for individual components to operate at a higher clock frequency than typical CMOS devices.
In this paper, we study the full-system implications of running the processing core and caches at ultra-high frequencies. Specifically, we introduce a modeling framework and perform and initial evaluation of out-of-order and in-order cores at high clock frequencies associated with superconductor electronics and cryogenic semiconductor computing [4]. We also wish to substantiate how other components in the system affect the performance, and how these other components should be designed to keep up with the faster components.
To evaluate the performance of cryogenic semiconductor and superconductor computing technologies, we:
@@
    Use gem5 [5, 6] to implement models of in-order and out-of-order cores operating at high clock frequencies, typical of cryogenic semiconductor and superconductor environments.
@@
    Choose workloads representative of real-world applications like NPB [7], SPEC CPU2006 [8], and GAPBS [9] to evaluate the performance of these components. These workloads provide a comprehensive understanding of how different types of applications respond to the ultra-fast processing capabilities offered by cryogenic semiconductor and superconductor technologies.
@@
    Simulate these workloads in different combinations of the aforementioned components in a superconductor clock domain and in the cryogenic semiconductor clock domain, while keeping the board and memory at their room temperature frequencies.
@@
This is a preliminary study to understand the potential of using superconductor electronics for general-purpose computing. Our results indicate that superconductor electronics may not be well suited for general-purpose computing, as the cache bandwidth limits the performance. However, we believe that superconductor electronics can be used for some applications. We conclude future work should focus on designing specialized accelerators that can take advantage of the potential speedups offered by superconductor electronics, instead of using them for general-purpose computing.
@@ II Background
In recent years, the pursuit of novel computing paradigms has led to a resurgence of interest in cryogenic semiconductor components and superconductor circuits. These technologies promise unprecedented speedups by exploiting the unique properties of materials at extremely low temperatures.
While superconducting was first discovered in 1911 [10], the theoretical concepts behind it were not fully understood until the BCS theory of the 1950s [11]. Due to further work in concepts like the Josephson effect [12] and flux quantization [13], superconductor circuits have been understood to operate at high clock frequencies of up to 100 GHz [14]. Therefore, superconductor circuits enable ultra-fast signal processing with minimal energy dissipation [3].
Cryogenic semiconductor components offer improved performance and energy efficiency by the reduced resistance at ultra-low temperatures. At cryogenic temperatures, semiconductor devices exhibit lower leakage currents and improved electron mobility, leading to improved performance. Therefore, these components also operate at a higher clock frequency, usually around 4 GHz [1].
Cryogenic semiconductor computing has gained significant attention in recent years, with researchers exploring cryogenic semiconductor components to enhance computing performance. Byun et al. introduced CryoCore, a cryogenic core based on the BOOM core, operating at 4.0 GHz with improved performance at cryogenic temperatures [1]. Min et al. proposed CryoCache, a cryogenic cache hierarchy leveraging these temperatures to enhance cache performance [15]. These components operate at ultra-low temperatures and offer substantial speedups compared to conventional semiconductor devices.
Our work represents a crucial step towards the practical implementation of superconductor and cryogenic semiconductor components in computing systems. Notably, we:
@@
    Lay the foundation for gem5 as a modeling framework for this field of computing research, providing all the tools necessary for evaluating the performance of computing systems under cryogenic and superconductor conditions.
@@ III Methodology
For all the experiments, we used gem5 [6, 5] as the simulator. We leverage the easy modification of gem5 CPU and cache models to make our own custom components, which is essential due to the microarchitectural details of the cryogenic semiconductor computing models available in the literature. We can also place the CPU, caches, and memory in different clock domains, which is essential for simulating cryogenic semiconductor computing environments. The gem5 Resources artifacts [16] provides workloads that we can also use for our experiments. Hence, gem5 is the ideal choice for our experiments.
@@ III-A Core Microarchitecture
We modeled the cryogenic semiconductor core and caches by building a model based on CryoCore by Byun et al. [1]. Their cryogenic core was based on BOOM [17], an out-of-order RISC-V core, and could run at 4.0 GHz. We used microarchitectural details provided in their paper to create a variant of the O3CPU in gem5. These microarchitectural details are enumerated in Table 3. For the pipeline values that were not specified, we went with the default values provided by gem5 for those variables. With regard to the latencies of the functional units and the branch predictor, we decided to go with the ones provided as part of the RISCVMatched prebuilt board [18] (based on the HiFive Unmatched and having a similar architecture to Rocket and BOOM [19]), since these values have been partially validated [20].
@@ III-D Experimental Setup
  • CryoCore and CryoCache (CryoAll) - Out-of-order CryoCore and CryoCache models in the cryogenic semiconductor clock domain (4 GHz).
  • SuperCore and CryoCache (SuperCryo) - CryoCore model in the superconductor clock domain (100 GHz) and CryoCache model in the cryogenic semiconductor clock domain (4 GHz).
  • SuperCore and SuperCache (SuperAll) - CryoCore and CryoCache models in the superconductor clock domain (100 GHz).
  • In-Order CryoCore and CryoCache (In-Order CryoAll) - In-order CryoCore model and CryoCache model in the cryogenic semiconductor clock domain (4 GHz).
  • In-Order SuperCore and CryoCache (In-Order SuperCryo) - In-order CryoCore model in the superconductor clock domain (100 GHz) and CryoCache model in the cryogenic semiconductor clock domain (4 GHz).
  • In-Order SuperCore and SuperCache (In-Order SuperAll) - In-order CryoCore and CryoCache models in the superconductor clock domain (100 GHz).
@@ IV Results
Research Questions 1. How does the performance of workloads vary with increasing clock frequency when subsystems are placed in a superconductor environment? 2. How is the performance of workloads affected by whether the core is out-of-order or in-order? 3. What new constraints must subsystems meet to leverage the improved performance resulting from increased clock frequency in a superconductor environment?
IV-A Workload Performance with Increasing Clock Frequency in Superconductor Environments
@@
Figure 2(a) and 2(b) show the speedup of the small workloads with respect to the out-of-order and in-order cores. Specifically, it compares the speedups achieved by placing the core and/or cache in a superconductor environment, for both out-of-order and in-order cores.
@@
At a high level, the takeaways from these results are that placing both core and cache in a superconductor environment significantly boosts performance, and that speedups of in-order superconductor components over their cryogenic counterparts are generally higher than the speedups of out-of-order superconductor components over their cryogenic counterparts. By characterizing the workloads that benefit the most from superconductor components, we also provide insights into the design trade-offs associated with these novel technologies.
@@
For out-of-order cores, Figure 2(a) shows the speedup of the small workloads in various out-of-order core configurations, with respect to the out-of-order CryoCore and CryoCache configuration. The bars represent different configurations for different small workloads. If only the core is placed in the superconductor environment, the performance improvement is not substantial, with the maximum speedup for small workloads being 2.7$\times$ for the cg workload over the baseline for an out-of-order core. Overall, the speedups for the small workloads range from 1.8$\times$ to 2.7$\times$ for the out-of-order SuperCore and CryoCache configuration, compared to the out-of-order CryoCore and CryoCache configuration. If both the core and the cache are placed in the superconductor environment, the performance improvement is more substantial for the out-of-order core. The maximum speedup for the small workloads for an out-of-order SuperCore and SuperCache configuration is 23.2$\times$ for the bt workload over the baseline out-of-order CryoCore and CryoCache configuration. Some workloads like lu and cg also receive speedups of more than 20$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration. However, some workloads like bfs and minisat do not receive substantial speedups, with the speedup being around 9.4$\times$ and 7.8$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration, respectively. Overall, the speedups for the small workloads range from 7.8$\times$ to 23.2$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration.
@@
Figure 2(c) shows the speedup of the large workloads in various out-of-order core configurations, with respect to the out-of-order CryoCore and CryoCache configuration. The bars represent different configurations for different SPEC2006 workloads. Similar to the small workloads, if only the core is placed in the superconductor environment, the performance improvement is not substantial, with the maximum speedup being 3.9$\times$ for the 400.perlbench workload over the baseline for an out-of-order core. Overall, the speedups for the large workloads range from 1.01$\times$ to 3.9$\times$, barring 456.hmmer, for the out-of-order SuperCore and CryoCache configuration, compared to the out-of-order CryoCore and CryoCache configuration. If both the core and the cache are placed in the superconductor environment, the performance improvement is more substantial for the out-of-order core. The maximum speedup for the large workloads for an out-of-order SuperCore and SuperCache configuration is 24.0$\times$ for the 456.hmmer workload over the baseline out-of-order CryoCore and CryoCache configuration. Some workloads like 435.gromacs and 401.bzip2 also receive speedups of more than 20$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration. However, most workloads do not receive substantial speedups, with the speedup being around 1.2$\times$ and 1.7$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration, respectively. Overall, the speedups for the large workloads range from 1.18$\times$ to 24.0$\times$ for the out-of-order SuperCore and SuperCache configuration, compared to the out-of-order CryoCore and CryoCache configuration.
@@
For in-order cores, Figure 2(b) shows the speedup of the workloads in various in-order core configurations, with respect to the in-order CryoCore and CryoCache configuration. The bars represent different configurations for different small workloads. If only the core is placed in the superconductor environment, the performance improvement is not substantial, with the maximum speedup for the small workloads being 2.3$\times$ for the tc workload over the baseline for an in-order core. Overall, the speedups for the small workloads range from 1.4$\times$ to 2.3$\times$ for the in-order SuperCore and CryoCache configuration, compared to the in-order CryoCore and CryoCache configuration. If both the core and the cache are placed in the superconductor environment, the performance improvement is more substantial for both the in-order core. The maximum speedup for an in-order SuperCore and SuperCache configuration is 24.3$\times$ for the bt workload over the baseline in the in-order SuperCore and CryoCache configuration. Some workloads like lu and cg also receive speedups of more than 20$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration. However, some workloads like bfs and minisat do not receive substantial speedups, with the speedup being around 8.4$\times$ and 9.6$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration, respectively. Overall, the speedups for the small workloads range from 8.4$\times$ to 24.3$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration.
@@
Figure 2(d) shows the speedup of the large workloads in various in-order core configurations, with respect to the in-order CryoCore and CryoCache configuration. The bars represent different configurations for different SPEC2006 workloads. Similar to the small workloads, if only the core is placed in the superconductor environment, the performance improvement is not substantial, with the maximum speedup being 2.6$\times$ for the 433.milc workload over the baseline for an in-order core. Overall, the speedups for the large workloads range from 1.1$\times$ to 2.6$\times$ for the in-order SuperCore and CryoCache configuration, compared to the in-order CryoCore and CryoCache configuration. If both the core and the cache are placed in the superconductor environment, the performance improvement is more substantial for both the in-order core. The maximum speedup for the large workloads for an in-order SuperCore and SuperCache configuration is 24.5$\times$ for the 456.hmmer workload over the baseline in the in-order SuperCore and CryoCache configuration. Some workloads like 435.gromacs, 444.namd and 401.bzip2 also receive speedups of more than 20$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration. However, most workloads do not receive substantial speedups, with the speedup being around 1.4$\times$ and 2.6$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration, respectively. Overall, the speedups for the large workloads range from 1.4$\times$ to 24.5$\times$ for the in-order SuperCore and SuperCache configuration, compared to the in-order CryoCore and CryoCache configuration.
@@
Therefore, the speedups vary significantly across different workloads, with similar workloads achieving similar speedups across different core architectures. While the speedups are consistent for smaller workloads, the variation in speedup is much wider when running real-world workloads. This variation suggests that even if we could build a conventional architecture in a superconductor environment, it is unlikely to benefit general-purpose workloads uniformly. Therefore, we need to either develop a new architecture specifically optimized for superconductor environments or apply superconductor technology to specialized accelerators where there is a small amount of memory traffic instead of general-purpose compute, which could leverage the low latency and high throughput benefits of superconductors more effectively.
@@
Takeaways The performance of workloads improved by placing components in the superconductor environment. The performance improvement was more substantial when both the core and the cache were in the superconductor environment. In-order cores were able to achieve a higher speedup than out-of-order cores when both the core and the cache were placed in the superconductor environment. The workloads that did not receive substantial speedups were bottlenecked by memory accesses. Therefore, memory systems would need to be redesigned to keep up with the faster core. The speedups vary significantly across different workloads, with similar workloads achieving similar speedups across different core architectures. While the speedups are consistent for smaller workloads, the variation in speedup is much wider when running real-world workloads. Therefore, we need to either develop a new architecture specifically optimized for superconductor environments or apply superconductor technology to specialized accelerators where there is a small amount of memory traffic instead of general-purpose compute.
@@
We compared the speedups of various configurations of out-of-order and in-order superconductor and cryogenic semiconductor cores and caches, with respect to the out-of-order CryoCore and CryoCache configuration. At a high level, we observed that the speedups are more substantial for out-of-order configurations compared to in-order configurations.
@@
From these figures, we can take away that for small workloads, placing components in a superconductor environment improves performance for both out-of-order and in-order cores. The performance improvement is more substantial when both the core and cache are in the superconductor environment. Interestingly, in-order cores achieve higher speedups compared to out-of-order cores when both components are superconductor. This could be because in-order cores are simpler and have less instruction-level parallelism, which could benefit from the higher clock frequency.
The out-of-order SuperCore and CryoCache configuration performs better than or equal to both the in-order CryoCore and CryoCache configuration and the in-order SuperCore and SuperCache configuration, for both the small and large workloads. Out-of-order cores can schedule instructions dynamically, which could benefit more from the higher clock frequency compared to in-order cores, which have a fixed schedule of instructions. Thus, even when all components are not superconductor, the out-of-order cores can still achieve significant performance benefits compared to in-order cores.